# Annotating Ambiguous Images: General Annotation Strategy for Image Classification with Real-World Biomedical Validation on Vertebral Fracture Diagnosis

### v1

@@ Abstract
While numerous methods exist to solve classification problems within curated datasets, these solutions often fall short in biomedical applications due to the biased or ambiguous nature of the data. These difficulties are particularly evident when inferring height reduction from vertebral data, a key component of the clinically-recognized Genant score. Although strategies such as semi-supervised learning, proposal usage, and class blending may provide some resolution, a clear and superior solution remains elusive. This paper introduces a flowchart of general strategy to address these issues. We demonstrate the application of this strategy by constructing a vertebral fracture dataset with over 300,000 annotations. This work facilitates the transition of the classification problem into clinically meaningful scores and enriches our understanding of vertebral height reduction.
Keywords:
Biomedical application annotation strategydata-centric Vertebral fracture
@@ 1 Introduction
The diagnosis of vertebral fractures is a critical field in the biomedical community. [15] states that the relative risk of death after a vertebral fracture is almost 9 times higher and that almost 20% of women will experience another fracture within 1 year after a vertebral fracture. The classification of vertebral fractures is primarily based on the height reduction of the vertebrae. Due to degenerative deformities, it is not easy to determine fractures on the height loss alone. This issue leads to inconsistent / ambiguous annotations of the data, which leads to inferior training of a neural network. In this paper, we present a strategy to produce high quality data for this task, independent of the network architecture used.
Our primary focus is on the application of vertebral fracture diagnosis. However, we will also present a general strategy for annotating and generating high-quality data for ambiguous classification tasks. Ultimately, we will circle back to the vertebral fracture diagnosis application to validate the effectiveness of this general approach.
In general, deep learning based methods yield state-of-the-art results in image classification [17]. However, their success is founded in large amounts of high quality data [3, 61]. While the amount of labeled data may be reduced with self- or semi-supervision for image classification, some high quality labeled data is still needed especially for domain-specific tasks [32]. However, a key challenge in annotating data is the reliability and consistency of annotators and thus their annotations. Annotations are typically created by humans and as such they may have different opinions or interpretations for difficult image classification tasks or when the images are ambiguous in nature. The literature concludes that a single annotation is not sufficient to the ensure high quality of labels in these cases [7, 13, 1, 50]. To address this challenge, soft labels as an average across multiple annotations have been proposed as a possible solution to this problem [45].
Despite the benefits of multiple annotations per image, annotation is costly and time-consuming. Thus, this approach is often not feasible for large datasets or even for smaller datasets where high-quality domain experts are required [25]. [44] proposes the use of proposal-guided annotations to overcome this issue. Proposal-guided annotation means that a preliminary trained network provides a rough estimate of the class for each image, which serves the annotators as guidance to the annotators in their decision making [8]. This guidance can help reduce the annotation time and increase the consistency and thus the overall quality [43, 46]
All of these challenges, low data quality and difficult labeling apply to vertebral fractures, but there are no well-established guidelines for precessing the data. Moreover, medical experts are in high-demand and cannot be burdened with annotating thousands of images multiple times. Our work makes several contributions to the field of data-centric deep learning for the application of vertebral fracture diagnosis and ambiguous image classification in general.
Our first core contributions are that we propose to unify the individual learning’s from the literature to a practical guide of how to annotate data, especially ambiguous data. Although previous work has explored the use of individual techniques,we provide a step-by-step strategy that covers all aspects of the annotation process. The goal of this guide is applicable to any image classification task to ensure high quality data.
Our second contribution is we verify the effectiveness of the proposed strategy by applying in on a dataset of vertebral fracture classification [35, 49]. The guidelines were defined before they were tested of the dataset, thus we can reflect and discuss the benefits of the proposed guidelines.
Importantly, we note that this is an application paper that aims to reflect and improve upon the existing literature, rather than introduce a novel method. By emphasizing the practical applications of our approach, we hope to facilitate the adoption of da ata-centric perspective in more research and applications, while contributing to the ongoing conversation about best practices in deep learning research.
@@ 2 Practical Guidelines: How to annotate ambiguous data?
Before outlining our practical approach, we will address the rationale for its necessity and explain the concept of ambiguous data to which these guidelines should be applied. Common labeling guidelines [40, 5] use hard-encoded labels (one correct class) and do not consider the issue that arises from the fact that people may disagree about the class for a given image. This can lead to the presence of ambiguous labels or, more generally, ambiguous data – an issue that is common in many real-world applications [52, 42, 23, 4, 22, 62]. The literature provides several reasons why one annotation per image is insufficient to capture the ambiguity of the data [54, 45, 44, 7, 13, 1, 50]. The hard label probability distribution for an image $x$ can be described as $P(L^{x}=\cdot)\in\{0,1\}^{K}$ with $K$ classes.
@@
Davani et al. showed that these multiple annotations should then be averaged to generate a soft label as in [45, 19], as opposed to using a majority vote, which is not optimal for ambiguous cases [7]. In addition, label smoothing, which involves smoothing hard-coded labels with a constant factor, is an established method for improving the performance of networks [55, 26, 36]. A soft-label probability distribution for $x$ can be described as $P(L^{x}=\cdot)\in[0,1]^{K}$.
It is important to note that ambiguous data is not just an error [39], but a data property. The data is inherently uncertain or ambiguous and thus must be treated as such. Thus, in general, the soft ground truth distribution $P(L^{x}=\cdot)$ is unknown and can only be approximated for the verification of the predicted results. This research is in line with a growing movement of data-centric deep learning [34, 21, 60, 13], where the focus is on data rather than model building.
@@
We give an overview about the practical guidelines in a flowchart in Figure 1. It is divided into five main steps: Defining the specifics of what, who, and how the annotation process will be performed, the process itself, and its post-processing.
@@
Figure 1: Flowchart with guidelines how to annotate ambiguous data, best viewed in color.
2.1 Definition - What?
The first step is to precisely define the classification task. In particular, the $K$ classes must be defined for the given problem and data. Additionally, raw unlabeled data must be collected, downloaded or generated based on the use case. The complete unlabeled image data is called $X_{r}$.
The next step is to determine the number of images to label. We looked at use cases with several thousand to several tens of thousands of images because we believe this is a common use case in real-world applications such as  [4, 43, 56, 10, 41]. When hundreds of thousands or even millions of images are available, self-supervised techniques may be applicable [6, 12]. However, even in these cases, performance evaluation requires a smaller labeled dataset that can be labeled with our strategy. We define $X_{u}\subseteq X_{r}$ as the subset containing only the samples that should be labeled.
The research we focus on builds on common supervised or semi-supervised deep learning [51, 53, 29]. Thus, we need enough samples per class to be available after annotation. The formulation ”enough samples” depends on the use cases, but our recommendation based on previously reported results would be that each class represents at least 1% of the complete data and that there are several dozen images per class, so that enough samples are available for evaluation [45]. If too few images are available, additional considerations about few or zero-shot learning approaches have to be taken into account [58, 57, 33]. However, this is beyond the scope of this paper.
For each image $x$ in a small labeled dataset $x\in X_{l}\subseteq X_{u}$ at least one hard-encoded ground truth label is needed. This data is used for annotator evaluation (see subsection 2.2) as well as potential training data of a proposal generation network (see subsection 2.3). An initial size of about 20% of the original dataset size was beneficial in [44], but it can be smaller or larger depending on the use-case.
2.2 Definition - Who?
A central question for the annotation process is who will annotate the image. Various concepts have been suggested in the literature, including the use of a single user, the collaboration of several workers, or even crowd-sourcing [40]. Depending on the background of the annotators, they need to be trained or selected for the given task [31]. To evaluate the quality of annoations, two elements are required in addition to the small labeled dataset for evaluation $X_{l}$: a threshold of acceptable performance $\mu$ and a suitable annotation platform.
The threshold $\mu$ depends on the evaluation metrics used. Macro accuracy or F1-score are commonly used and can show a high correlation [28, 16, 2, 45, 24, 30]. The results in [46] indicate that a threshold $\mu$ of 60% to 80% is often reasonable enough if the annotations are aggregated in the end. When choosing the threshold, one should keep in mind that the quality of the labels in $X_{l}$ affects the maximum achievable scores. For example, due to intra- or inter-observer variability, a score higher than 80% may not be possible [47].
Based on the task definition in subsection 2.1, a few classes with many occurrences are expected, so Sagar et al. recommend using a browsing environment for the annotation process, where multiple images can be viewed side by side [40]. However, other annotation platforms such as [48] are possible. These platforms may introduce an additional overhead during the setup, but can dramatically increase the speed of annotation.
The annotators should be trained until they exceed the desired performance $\mu$ on the labeled data $X_{l}$ with the chosen platform.
2.3 Definition - How?
The key question in this section is whether or not to use proposals in the annotation process. This question is central due to the fact that proposals can reduce the cost of annotation by a multitude, but may introduce unwanted side effects [20, 44]. This side-effect / bias can be introduced by to the fact that annotators are more likely to accept a proposed class, resulting in a skewed label distribution $P(L^{x}=\cdot)$ towards the proposed class. For this decision process, several parameters need to be approximated or at least their way of approximating them later should be defined. These parameters are motivated by [44] and are the class confusion matrix $c$ and the expected speedup $S$. We will first describe how these variables are used and refer to the supplement for possible approximation techniques. The most important part is to understand, when specific variables are needed, how they could be approximated at this step or how they will be approximated if they are only needed later.
The decision of whether to use proposals is guided by two main questions from [44]. The first question concerns the acceptability of introducing a bias. Although introducing a bias may seem undesirable, it is a bias that has been approved by human annotators. In recommendation systems [8, 9, 14], this bias is often inherently accepted, and even when humans reach a consensus [35], they may influence each other. If the conditions allow for an acceptable bias, proposals should be utilized. If not, the expected speedup must be evaluated.
The benefit of proposals is greater with a higher speedup ($S$), as a higher speedup enables easier reversal of the introduced bias with more annotations. However, there is a trade-off point where the speedup is too small to justify the introduced bias, and it becomes unlikely that the bias can be recovered with an equal or smaller budget. In [44], this tradeoff point is defined to be around 3. If the speedup falls below this threshold, the recommendation is to avoid using proposals and resort to normal annotations. Conversely, if the speedup exceeds the threshold, proposals can be employed.
The class confusion matrix $c$ is used in post-processing (see subsection 2.5) to improve the final distribution and is not needed until then.
2.4 Annotation Process
This step involves the most work because it concentrates on creating the annotations. Several preparatory steps need to be taken, depending on whether the utilization of proposals is intended or not. Previous literature has shown that proposals improve the consistency of annotations [38, 8], especially the use of overclustering seems to be beneficial with ambiguous data [43]. Therefore, we recommend using DC3 proposals [46], which have been shown to perform better on real-world ambiguous data than common network predictions and work with many semi-supervised methods [51, 30, 53]. The initial labeled data $X_{l}$ (defined in subsection 2.1) can be used as training data for such an arbitrary neural network. The network should be trained until it reaches the same quality threshold $\mu$ as the annotators. Before starting the annotation process, two variables must be defined based on the decision in subsection 2.3: the number of annotations until consensus ($A_{c}$) and the total number of annotations ($A^{c}$). These variables play a crucial role in determining the trade-off point between achieving the best approximation of the underlying distribution $P(L^{x}=\cdot)$ and minimizing the annotation effort and associated costs.
$A^{c}$ represents the general number of annotations per image, reflecting the quality of the approximation. While some data in $X_{u}$ may exhibit ambiguity and result in disagreement among annotators, there is a portion of data where all annotators reach a consensus. Although the specific proportion of complete consensus ($p_{c}$) has not been reported, [46] reports the proportion of at least 95% consistent annotations ($\hat{p_{c}}$) ranging from about 50% to 80% across four different datasets. To reduce the total annotation costs, we propose annotating only those images that do not achieve complete consensus after $A_{c}$ annotations. This approach aims to identify images where complete consensus can be achieved with fewer annotations than $A^{c}$.
The values chosen for $A_{c}$ and $A^{c}$ depend on the specific classification problem and previous design choices. In the supplementary section, we discuss plausible numbers of annotations per hour ($\frac{a}{h}$) and potential 95% confidence guarantees for different combinations of $A_{c}$ and $A^{c}$.
For the case where $A_{c}=10$ and $A^{c}=50$, two main conclusions can be drawn. First, based on the previously reported $\frac{a}{h}$ in [46], it is likely that all annotations can be completed within 10 days by a single annotator. This workload seems manageable, even for smaller projects, especially when shared among multiple annotators. Second, considering the approximation of $P(L^{x}=\cdot)$ and its unknown true value, if we want to achieve an approximation within a margin of 0.15 around the unknown true value with at least 95% confidence, using $A_{c}=10$ and $A^{c}=50$ is a reasonable choice.
2.5 Post-Processing
As discussed in subsection 2.3, the use of proposals during the annotation process introduces a bias due to the default effect [20]. It is important to note that while this speeds up the annotation process and produces more consistent annotations, it may also introduce a bias that may not be acceptable. Schmarje et al. propose in [44] with CleverLabel a combination of classblending and bias correction to improve the approximated distribution, which should be applied if proposals were used during the annotation process. If no proposals were used to create the annotations, it is still recommended to improve the distribution $P(L^{x}=\cdot)$ by blending with the class distribution $c$.
@@ 3 Evaluation
The goal of this paper is to solve real-world application annotation tasks such as the medical task of vertrebral fracture classification. For this purpose, we will describe how the previously defined strategy can be applied to this task on the publicly available datasets Verse2019[35] and Verse2020[49]. An analysis of the benefits or shortfalls of the decisions is given in subsection 3.2.
@@ 3.1 Applying the Strategy
We use the Verse datasets for better reproducibility of our osteoporotic vertrebral fracture classification. As in [35], we consider only at thoracolumbar vertebrae because osteoporotic fractures occur mainly in this region of the spine, as shown in Figure 2. For the classification of the fractures, we use an adaptation of the most commonly used score, the semi-quantitative Genant score [11]. As shown in Figure 2, the score defines four classes, which are mainly guided by the height reduction of the vertebrae compared to its neighbors or unfractured vertebrae. However, degenerative deformities (e.g. short vertebral height) of up to 20% are ignored for the osteoporotic fracture classification. Partially, these degenerative cases can not be assessed in the dataset because the resolution is lower than in normal radiographs. As shown in Table 1, this results in 3,761 individual vertebrae with a high class imbalance towards class zero. Overall, we have at least 1.1% samples per class and can apply common semi- or supervised techniques.
@@
[…]
Figure 3: Illustration of different image modalities
Unlike to the original author of the Verse2019 [35] dataset, we will not use medical experts for the annotation and will use 2D projections of the original 3D CT data.
Our decision to use non-medical experts has two main reasons. First, we want to show that non-experts can successfully annotate data with the help of proposals. Second, we plan to generate tens of thousands of annotations, which would not be feasible with medical experts, who are in high demand. While we will not use medical experts for the majority of the annotations, we have worked closely with the radiologist at XXXX (censored for review). We discussed our design choices with them, particularly regarding the definition of the task and the representation of the vertebrae used. As mentioned above, we trained the hired workers on this specific task and ensured that they achieved a macro accuracy and F1 score of at least $\mu=0.6$ on the complete dataset with and without proposals for at least two scores before they could annotate the final scores.
We do not use the original 3D CT data (see Figure 3) for the annotation but rather a 2D projection of the central slices for each vertebrae (see Figure 3) and its segmentation (see Figure 3). Additionally, we show the vertebrae above and below the target vertebrae to allow an evaluation of the height reduction. A blended example is shown in Figure 3. There are three major benefits/reasons of using 2D projections over 3D CT data. First, the Genant scores were defined on 2D X-Ray and 2D projections of CT data are visually very similar to X-Ray and thus are better comparable. Second, we want to use the reported methods from the literature for the proposal generation [46], which were only tested on 2D data. While it should be theoretically possible to extend the used networks to 3D, such major structural changes could lead to an inferior result which would dilute the results, such experiments may be future work. Lastly, when it comes to diagnostics, radiologists mainly look at one central slice which we include and the major 3D structures can be detected on this projection.
We hired 5 workers to annotate vertebrae in a web-based environment developed for this evaluation. Public annotation platforms do not support the combination of clusters and classes that we want to evaluate in this paper. Details about the platform are given in the supplementary. As $X_{l}$ we used the partially available labels from [35].
In a medical consensus process in an expert panel, the proposals of different people are unified, but influences between people due to social dynamics cannot be excluded. Therefore, we argue that a bias that may be introduced due to human approved annotations is comparable to a consensus process and therefore acceptable. This means that we will use proposals during the annotation and $c$ and $S$ only need to be approximated after the annotation process based on the defined strategy. To allow a validation of our decisions, we will also raise annotations without proposals and thus we can approximate the class distributions $c$ and speedup $S$ based on these annotations. Additional approaches to approximate $S$ as well as the approximation of a data-specific offset $\delta$ (see [44]) are discussed in the supplement.
Based on the training annotations, we estimated the $\frac{a}{h}$ to be about 2,000 and 3,000 without and with proposals, respectively. These results are similar values previously reported in [46]. For $A_{c}=10$, $A^{c}=50$ and estimated $p_{c}=0.5$, this would result in about 250,000 annotations with and without proposals. Furthermore, based on the estimated $\frac{a}{h}$ of about 2,500, this resulted in a workload of about 20 hours per annotator which was within our given budget. So we chose the number of annotations at the recommended values of ${A_{c}=10,~{}A^{c}=50}$. Additionally, we planned to annotate the 3D data to evaluate the quality difference between 2D and 3D. We limited these annotations to the test set of the original Verse2019 dataset [35].
3.2 Analysis
The annotators completed more than 100 iterations of annotation, half of them with proposals and half without. However, one annotator did not meet the acceptance threshold $\mu$ and was excluded along with the training iterations, leaving 80 valid iterations or a total of about 250,000 annotations. Additionally, 11 iterations were performed using 3D data from the test dataset for comparison. The average annotation rate was calculated to be 3,259.36 $\frac{a}{h}$, which means that it takes about one hour to complete the 3,761 annotations with the additional technical overhead and pauses. The comparison of the resulting dataset-specific variables with previously reported values [46] is given in Table 1. We see that the dataset offset $\delta$ and the rate of near-consensus rate $\hat{p_{c}}$ are similar to previously reported results. It is important to note that $p_{c}=0.27$, which is significantly smaller, resulting in more data with $A^{c}$ number of annotations. It should be evaluated in the future, if instead of total consensus a high number of agreeing votes is sufficient to reduce the cost of the annotation process. The largest class is greater than in the previous reports and the speedup $S$ is smaller in Table 1. This relationship is theoretically justified because a proposal is less important since the majority of images already belongs to a class and thus defines some kind of a proposal.

### v2

# Annotating Ambiguous Images: General Annotation Strategy for High-Quality Data with Real-World Biomedical Validation

@@ Abstract
In the field of image classification, existing methods often struggle with biased or ambiguous data, a prevalent issue in real-world scenarios. Current strategies, including semi-supervised learning and class blending, offer partial solutions but lack a definitive resolution. Addressing this gap, our paper introduces a novel strategy for generating high-quality labels in challenging datasets. Central to our approach is a clearly designed flowchart, based on a broad literature review, which enables the creation of reliable labels. We validate our methodology through a rigorous real-world test case in the biomedical field, specifically in deducing height reduction from vertebral imaging. Our empirical study, leveraging over 250,000 annotations, demonstrates the effectiveness of our strategies decisions compared to their alternatives.
Keywords: Data quality management, ambiguous data, annotation data application, data-centric AI, vertebral fracture
@@ 1 Introduction
Deep learning methods are at the leading methods for image classification, dependent on the availability of substantial high-quality data (Beyer et al., 2020; Yun et al., 2021). While techniques like self- or semi-supervision can reduce the need for labeled data, high-quality labels remain a necessity, especially in domain-specific tasks (Liu et al., 2023).
A critical challenge lies in the reliability and consistency of human annotations, especially for ambiguous or complex classification tasks. The consensus in current research suggests that single annotations are inadequate for ensuring label quality in such cases (Davani et al., 2022; Grossmann et al., 2022; Basile et al., 2021; Sharmanska et al., 2016). To mitigate this, the adoption of soft labels, derived from averaging multiple annotations, has been proposed (Schmarje et al., 2022a). Soft labels can capture the inherent data uncertainty, which is different from the model uncertainty often assumed by uncertainty estimation methods. However, obtaining multiple annotations per image is resource-intensive and impractical for large datasets or when expertise is scarce (Krizhevsky et al., 2012). Schmarje et al. (2023) introduces proposal-guided annotations, where a pre-trained network provides preliminary class estimates to guide annotators, thereby enhancing annotation efficiency and quality (Desmond et al., 2021; Schmarje et al., 2021a, 2022b). An illustrative example of this process is shown in Figure 1.
In this paper, we present a comprehensive strategy for generating high-quality data in ambiguous classification scenarios. High quality data means that we improve the label quality in comparison to one annotation only. We validate our approach in vertebral fracture diagnosis, a field with significant challenges due to data ambiguity, which is ideal for testing our strategy’s efficacy.
Figure 1: Illustration of the concept of hard and soft labels and how they can be created from annotations – The recommended process has three steps. In the first step, an image $x$ is selected for annotation. The unknown ground-truth distribution ($P(\hat{L}^{x}=\cdot)$) could either be soft or hard as shown by the examples in the lower half. During the annotation, multiple annotations are created either with or without proposal. A proposal means that one class is recommended during the annotation process. In the example, class B is proposed and 32 annotations are generated. The average across these annotations could already be used as an approximation of the soft-label $P(L^{x}=\cdot)$, however it might be biased towards the proposal since it is more likely to accept a proposal Schmarje et al. (2023). Our post-processing step enhances the approximated distribution ($P(L^{x}=\cdot)$)from the second step by reducing this bias. In the provided example, the probability of class B is reduced since it was most likely overestimated due to the used proposal of class B.
The diagnosis of vertebral fractures holds critical importance in medicine. As noted by Haczynski and Jakimiuk (2001), vertebral fractures significantly increase mortality risks and recurrence rates. Classifying these fractures, primarily based on vertebral height reduction, is challenging due to degenerative changes, leading to inconsistent annotations and impacting neural network training. This scenario underscores the necessity of our approach, especially when domain experts are scarce.
Our primary contribution lies in synthesizing literature insights into a practical, step-by-step annotation guide for ambiguous real-world data. While individual annotation techniques have been explored previously, our unified strategy addresses the entire annotation process comprehensively. We aim to make this approach universally applicable to various image classification tasks, ensuring high-quality data. Furthermore, we will release software guidelines to assist in creating high-quality annotations following our approach.
Our second key contribution is empirically validating our strategy on a vertebral fracture dataset (Löffler et al., 2020; Sekuboyina et al., 2021). By defining the guidelines prior to dataset testing, we provide an objective evaluation, underscoring the real-world applicability of our method.
By focusing on practical applications, this work seeks to promote a data-centric perspective in research and application, contributing to the discourse on data quality best practices.
@@ 2 Practical Guidelines: How to Annotate Ambiguous Data?
Before detailing our approach, it’s essential to understand why these guidelines are crucial, especially in the context of ambiguous data. Typical labeling guidelines (Sager et al., 2021; Chang et al., 2017) rely on definitive, hard-encoded labels and often overlook the disagreement among annotators about a given image’s class. Such discrepancies lead to ambiguous labels or, in broader terms, ambiguous data – a prevalent issue in numerous real-world applications (Tarling et al., 2021; Sambyal et al., 2022; Karimi et al., 2020; Brünger et al., 2019; Jiang et al., 2021; Zhang et al., 2023). The probability distribution of a hard label for an image $x$ can be represented as $P(L^{x}=\cdot),P(\hat{L}^{x}=\cdot)\in\{0,1\}^{K}$, where $K$ is the number of classes and $L^{x},\hat{L}^{x}$ are random variables mapping an image $x$ to its class probability. We distinguish between $L^{x}$ and $\hat{L}^{x}$, which represent the estimated and ground-truth probability distributions for image $x$, respectively. The literature underlines that a single annotation per image is inadequate for capturing data ambiguity (Uma et al., 2021; Schmarje et al., 2022a, 2023; Davani et al., 2022; Grossmann et al., 2022; Basile et al., 2021; Sharmanska et al., 2016), necessitating multiple annotations for high-quality data.
@@
Davani et al. (2022) demonstrated that averaging multiple annotations to form a soft label, as proposed in (Schmarje et al., 2022a; Hemming et al., 2018), is superior to using a majority vote in ambiguous cases. Moreover, label smoothing, a technique that tempers hard-coded labels with a constant factor, has been established as a method for enhancing network performance (Vaswani et al., 2017; Krothapalli and Abbott, 2020; Lukasik et al., 2020). A soft-label probability distribution for image $x$ can be expressed as $P(L^{x}=\cdot),P(\hat{L}^{x}=\cdot)\in[0,1]^{K}$.
It is crucial to recognize that ambiguous data is not merely an error, as suggested in (Park and Chung, 2022), but a characteristic of the data itself. Such data inherently possess uncertainty and should be treated accordingly. As such, the soft ground truth distribution $P(\hat{L}^{x}=\cdot)$ is generally unknown and can only be approximated for validating predicted results. This approach aligns with the growing emphasis on data-centric deep learning (Liu et al., 2021b; Jarrahi et al., 2022; Whang et al., 2023; Grossmann et al., 2022), which prioritizes data quality over model architecture.
@@
We present a comprehensive overview of our practical guidelines through a flowchart in Figure 2. The guidelines encompass five main steps: defining the specifics of what, who, and how the annotation process will be executed, the annotation process itself, and its post-processing. The subsequent sections provide insights into the pipeline, with detailed discussions in Appendix A and Appendix B.
@@
Figure 2: Flowchart with guidelines on how to annotate ambiguous data, best viewed in color.
  1. 1.
    Definition - What?  This stage outlines the classification task, specifying the classes ($K$) and gathering raw, unlabeled image data ($X_{r}$). A crucial element is selecting a representative subset ($X_{u}$) for annotation. The selection of $X_{u}$ emphasizes the importance of having an adequate sample size per class for effective training and evaluation or considers the use of self-supervised techniques for datasets without enough representative information. The definition of $X_{l}$ describes a smaller, precisely labeled dataset for evaluating annotators and training proposal generation networks.
  2. 2.
    Definition - Who?
    This stage focuses on identifying suitable annotators for image labeling, ranging from individuals to crowdsourcing. It is important to train the annotators according to the task’s specific needs and setting a quality threshold for annotations ($\mu$), typically recommended between 60% and 80%. For classes with higher occurrence rates, a browsing environment for annotation is advised to improve process efficiency.
  3. 3.
    Definition - How?
    In this stage, the decision to include proposals in the annotation process is discussed. Proposals can accelerate annotation but may introduce bias as discussed in (Schmarje et al., 2023). Key considerations include the acceptability of bias, the extent of speedup ($S$)), and a class confusion matrix ($c$) for later processing. Proposals are recommended if the bias is manageable and the speedup significant (typically above a threshold of 3), otherwise, standard annotations are preferred.
  4. 4.
    Annotation Process
    This phase involves conducting the annotations, with or without proposals. For ambiguous data, overclustering (Schmarje et al., 2021b, a) is advised, and DC3 (Schmarje et al., 2022b) proposals are recommended for better performance. We propose to separate the annotations process into the number of annotations needed for early consensus ($A_{\text{cons}}$) and the total annotations for difficult cases($A^{\text{cons}}$). This separation balances ground-truth accuracy against the effort and cost of obtaining annotations.
  5. 5.
    Post-Processing
    The concluding phase addresses the bias potentially introduced by proposals in the annotation process. CleverLabel (Schmarje et al., 2023), combining class blending and bias correction, is proposed for refining the approximated distribution if proposals were used. Even without proposals, blending the distribution with the class distribution ($c$) is recommended to enhance it.
@@ 3 Evaluation
Our objective is to address the challenges of annotating real-world application tasks, such as classifying vertebral fractures in medical images. To demonstrate the practicality of our proposed strategy, we apply it to the task using publicly available datasets Verse2019(Löffler et al., 2020) and Verse2020(Sekuboyina et al., 2021).
@@ 3.1 Applying the Strategy
The Verse datasets were chosen for their reproducibility and suitability in validating our workflow for classifying osteoporotic vertebral fractures. Following Sekuboyina et al. (2021); Löffler et al. (2020), we focus on thoracolumbar vertebrae, the primary site for osteoporotic fractures, as illustrated in Figure 4. Our classification approach is based on an adapted version of the Genant semi-quantitative score (Genant et al., 1996), which categorizes fractures into four classes based on the degree of height reduction in vertebrae compared to their neighbors. However, we exclude degenerative deformities, such as minor height reductions (up to 20%), which are often indistinguishable in the given CT images due to lower resolution compared to standard radiographs. As detailed in Figure 4, this results in a dataset of 3,761 individual vertebrae, with a notable class imbalance skewed towards class zero.